### v1

@@ 4 RQ1: Popularity Trend Analysis
Fig. 3 illustrates the rising trend in OpenAI’s development popularity, as evidenced by the increasing number of posts on the OpenAI forum and the surge in registered users. The figure begins at the forum’s inception, tracking the growth in new posts and users every three months. Given that the user number significantly exceeds the post number, we apply a logarithmic transformation to the actual numbers for a clearer representation of the results. This visualization reveals a significant increase in developer interest in OpenAI since 2021, underscoring the timeliness and urgency of our empirical study.
@@
Figure 3: The number of new posts and users every three months since creating the OpenAI forum. Notice we apply a logarithmic transformation to the actual numbers for a clearer representation of the results.
Before November 2022, the growth trend was fairly consistent. However, the introduction of GPT-3.5 and its fine-tuned product, ChatGPT, to the public by OpenAI in late November 2022, marked a substantial shift. The number of posts increased by 292%, and the number of users rose by 1,302% compared to the preceding period. The release of GPT-4 to paying customers in March 2023 further accelerated this trend, with a 323% increase in posts and a 628% increase in users. This vigorous growth has persisted to the present.
@@
Finding 1. The introduction of GPT-3.5 and ChatGPT in November 2022, followed by GPT-4 in March 2023, significantly accelerated the growth of the OpenAI community, markedly increasing the number of posts and users on the OpenAI forum.
5 RQ2: Difficulty Level Analysis
Fig. 4 illustrates the proportion of questions with varying numbers of replies. The figure shows that 54% of questions receive fewer than three replies, while only 10% receive more than ten replies. This suggests that questions related to OpenAI development are generally challenging to resolve.
Figure 4: The proportion of questions with different numbers of replies.
@@ 5 RQ2: Difficulty Level Analysis
We summarize potential reasons as follows. First, the new and complex technological domains of OpenAI require deep expertise and skills from developers. This scarcity of qualified professionals results in limited responses. Second, the rapid evolution of OpenAI’s technology often leaves issues unresolved, leading to fewer available answers. Finally, the lack of comprehensive documentation and tailored support resources makes it difficult to address diverse developer needs, prolonging the resolution process for many questions.
@@
Finding 2. Problems related to OpenAI development often face challenges in receiving immediate and sufficient replies.
6 RQ3: Challenge Taxonomy Construction
@@
Fig. 5 illustrates the hierarchical taxonomy of challenges faced by OpenAI developers. In this figure, the nodes are shaded in descending levels of grey based on their depth in the hierarchy (e.g., leaf nodes are white). Each leaf node represents a specific category, while its parent node is an inner node composed of multiple subcategories. For example, API (B) is an inner category that can be further classified into four leaf categories: Faults in API (B.1), Error Messages in API Calling (B.2), and API Usage (B.3).
@@
The proportion of posts for each category is shown in parentheses. In summary, our taxonomy includes six inner categories and 27 leaf categories. Based on this taxonomy, we find that OpenAI developers encounter a wide variety of problems across different aspects of development, highlighting the diversity of challenges in OpenAI development. Next, we describe and provide examples for each category and summarize our findings and implications for developers and the OpenAI organization.
Figure 5: Our Constructed Challenge Taxonomy for OpenAI Developers.
Finding 3. The challenges faced by OpenAI developers are multifaceted and diverse, encompassing 27 distinct categories. These range from Conceptual Questions to API Usage, from Prompt Design to Text Generation, and from Rate Limitation to Regulation.
6.1 General Questions
@@
General questions encompass challenges that lack specific implementation details and are often posed by developers seeking foundational knowledge in OpenAI development. Our findings indicate that 26.3% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@
Integration with Custom Applications (A.1). With the launch of ChatGPT and the availability of OpenAI’s APIs, more developers are leveraging these GPTs/APIs to build custom applications, such as translation systems and text classification systems. When integrating the API into their applications, developers often face complex and diverse technical questions. For example, developers seek to understand how to integrate custom GPTs with web page interfaces [^8] https://community.openai.com/t/589076 or inquire about the feasibility of implementing specific functionalities through API calls, such as considering specialized terminology that varies across different languages [^9] https://community.openai.com/t/584762, as shown in Fig. 6. Additionally, some developers raise concerns about the reproducibility of results in commercial text classification systems using GPT-3.5-turbo, especially after model updates [^10] https://community.openai.com/t/281658. Our findings indicate that challenges in this subcategory account for a significant 17.0%, highlighting the diverse and complex technical issues developers face when building their custom applications.
@@
Figure 6: A sample post in the integration with custom application (A.1)
@@
Conceptional Questions (A.2). This category encompasses questions about basic concepts or background knowledge of OpenAI development, such as how to calculate the tokens needed for each question [^11] https://community.openai.com/t/81018, whether the GPT-4 API has access to the Internet [^12] https://community.openai.com/t/468615, or whether plugins can be used with the API [^13] https://community.openai.com/t/186333. Additionally, developers inquire about the internal implementation details, such as mechanisms of the content filter [^14] https://community.openai.com/t/1095. In our taxonomy, the posts in this subcategory account for 6.4%.
@@
Finding 4. The majority of the total challenges, 26.3%, fall under the category of General Questions. Within this category, the subcategory Integration with Custom Applications represents the largest proportion, accounting for 64.6% of General Questions.
@@
Discussion and Implication: General questions often reflect fundamental challenges developers face when working with OpenAI technologies. These challenges span various aspects (such as application integration, conceptual questions, and feature suggestions). When leveraging OpenAI’s APIs to build their applications (such as translation systems, text analysis systems, speech recognition systems, and others), developers often encounter complex technical questions. These include API integration methods, performance issues, output reproducibility issues, interpretability of output content, and so on. These challenges highlight the diverse and intricate nature of API integration and the need for clearer guidelines and examples from OpenAI to assist developers in these areas. As integration with custom applications is a major challenge, OpenAI could develop dedicated resources and support mechanisms to streamline this process. This may include detailed tutorials and integration best practices. Questions in the subcategory of conceptual questions involve basic concepts and background knowledge necessary for understanding OpenAI’s technologies. Developers often inquire about token calculations, API capabilities and internal mechanisms, model versions, and other related issues. These challenges underscore the need for improved foundational documentation. To mitigate these challenges, improving educational resources will help developers grasp fundamental concepts more effectively, reducing the frequency of basic questions and accelerating their development process. For the last subcategory, developers frequently suggest new features or improvements related to API operations, such as better cost management tools, enhanced plugin and GPTs recommendation features, and support for a wide range of data formats. This subcategory reflects a proactive interest from developers in enhancing OpenAI’s tools. Therefore, OpenAI should establish a structured feedback mechanism to capture and address feature suggestions from developers. This will help prioritize feature development based on actual user needs and enhance overall user satisfaction.
6.2 API
@@
This category contains challenges related to the OpenAI APIs. In our taxonomy, 22.9% of the total challenges fall under this category. The detailed subcategories are discussed as follows.
@@
Faults in API (B.1). When calling OpenAI APIs, developers frequently encounter a variety of issues, such as low-quality generated content, limitations in model comprehension, text coherence problems. These issues often result in outcomes that do not meet developers’ expectations. The majority of these issues are related to unsatisfactory output, such as the presence of extraneous information (like spaces and newlines) in the API’s responses [^17] https://community.openai.com/t/578701, as well as phrase repetition in answers [^18] https://community.openai.com/t/54737. Beyond these common faults, developers face other types of faults, such as APIs failing to respond due to requests being too frequent [^19] https://community.openai.com/t/483090. Additionally, when dealing with specific technical problems, such as converting natural language into SQL queries, they might produce incorrect answers [^20] https://community.openai.com/t/426293. These examples illustrate the diverse technical challenges that can arise. This subcategory represents 8.7% of the total challenges.
@@
Error Messages in API Calling (B.2). This category primarily includes various error messages that developers encounter when calling OpenAI APIs. For instance, developers might face request errors [^21] https://community.openai.com/t/154711 and data value capacity limit errors when calling an API for image editing [^22] https://community.openai.com/t/464513. As shown in Fig. 7, developers may also face issues related to the OpenAI server, such as gateway timeout errors [^23] https://community.openai.com/t/219369. This subcategory represents 7.5% of the total challenges.
Figure 7: A sample post in the error messages in API calling (B.2)
@@
API Usage (B.3). Parameter configuration is a principal concern among developers, necessitating the parameter value adjustment to meet specific requirements and generate expected outputs from the API. A significant portion of these problems center on determining the appropriate values for API parameters [^24] https://community.openai.com/t/477563 (as shown in Fig. 8) and understanding the internal working and implementation mechanism (such as the response_model parameter [^25] https://community.openai.com/t/511738). Additionally, some challenges are closely associated with GPT models. For example, setting identical values for the seed parameter does not result in consistent outputs for specific GPT models [^26] https://community.openai.com/t/487245. This subcategory represents 6.7% of the total challenges.
Figure 8: A sample post in the API usage (B.3)
@@
Discussion and Implication: Our results show that developers frequently encounter various issues when using OpenAI APIs. Some of these problems are related to the specific GPT models being used. For instance, certain APIs only support the gpt-4 turbo model for specific fields like response_format, indicating potential compatibility issues. Therefore, OpenAI should ensure that their documentation clearly outlines the compatibility of different models and fields, and provide detailed examples and best practices for developers to follow. This can help mitigate confusion and reduce the frequency of these issues. Developers also encounter various error messages when calling OpenAI API. To address these issues, OpenAI should consider providing more detailed error messages that help developers understand the cause of the errors and how to fix them. Additionally, offering guidance on common parameter configurations and their implications can be beneficial. For example, addressing phrase repetition issues can be managed by adjusting the frequency_penalty parameter, where setting positive values can penalize new tokens based on their existing frequency in the text so far, thereby decreasing the model’s likelihood of repeating the same line verbatim. Parameter configuration is a principal concern among developers. To help developers, OpenAI should provide comprehensive documentation that includes detailed explanations of parameter values and their effects. Offering sample configurations and use-case scenarios can guide developers in making informed decisions about parameter settings. For example, explaining how different values for the temperature parameter affect the creativity and randomness of the generated content can help developers tailor the API output to their needs. In summary, OpenAI should ensure that the API documentation is thorough, up-to-date, and includes detailed examples and best practices. Moreover, providing more informative error messages can help developers quickly identify and resolve issues. Finally, OpenAI should monitor community forums and address recurring issues by updating the API and documentation accordingly.
6.3 Generation and Understanding
@@
Due to powerful generation and understanding capabilities, developers can use OpenAI models to perform a variety of tasks (such as text processing, image processing, and audio processing) or fine-tune the OpenAI models for downstream tasks. the challenges in these tasks often overlap with other challenge categories, so a question (such as parameter setting issues related to image generation [^27] https://community.openai.com/t/359438) may be classified into multiple categories. Challenges related to this category account for 19.9% of the total challenges. The detailed subcategories are discussed as follows.
@@
Text Processing (C.1). As OpenAI’s core capability, text generation helps developers to create coherent and semantically sound textual content based on the input text (i.e., the prompt). This feature allows developers to automatically generate various types of text (such as articles, stories, and dialogues). By modifying the input prompts, developers can control the style, content, and length of the generated text to meet diverse application requirements across different scenarios. This category contains challenges associated with text generation, such as encountering errors when using the text generation API [^28] https://community.openai.com/t/486260 or getting repeated responses from the davinci model [^29] https://community.openai.com/t/254241. Moreover, developers want to provide feedback on the given output to improve the upcoming outputs when using the gpt-3.5-turbo model [^30] https://community.openai.com/t/452220. This subcategory constitutes 6.8% of all challenges.
@@
Fine-tuning GPT Models (C.2). OpenAI developers can fine-tune the models provided by OpenAI by using the training datasets of the downstream tasks. Due to the widespread use of fine-tuning, developers frequently face questions related to this functionality. For example, they often inquire about how to set the parameter n_epochs [^31] https://community.openai.com/t/339246 or request the fine-tuned model to output results in JSON format [^32] https://community.openai.com/t/525169. Additionally, developers are interested in whether they can iteratively train a fine-tuned model [^33] https://community.openai.com/t/18248. This subcategory of questions accounts for 6.7% of the total challenges.
@@
Image Processing (C.3). Developers can influence image generation by inputting prompts into the model, allowing for the creation of images with customized features and styles. This versatile feature is useful across diverse applications, including image editing and artistic creation, offering developers opportunities for creative expression and design. However, developers often encounter challenges such as request errors when seeking image generation solutions [^34] https://community.openai.com/t/586006 or perceiving that the quality of the generated images is lower than expected [^35] https://community.openai.com/t/435187. Additionally, they are concerned about whether the generated images can be used freely without violating intellectual property rights [^36] https://community.openai.com/t/23140. This subcategory constitutes 2.5% of all challenges.
@@
Embedding Generation (C.4). By using the Embedding APIs provided by OpenAI, developers can leverage powerful text representation capabilities to convert text into high-dimensional vector representations. These vectors can capture the semantics and contextual information of the text, which can be used for tasks such as semantic search and text clustering. For example, developers may encounter PermissionError when creating embeddings [^37] https://community.openai.com/t/587579. Additionally, they often inquire whether modifying user queries can improve the accuracy of semantic search [^38] https://community.openai.com/t/393047 and how to determine the length of embedded content [^39] https://community.openai.com/t/111471. Challenges related to this subcategory account for 1.8% of the total challenges.
@@
Audio Processing (C.5). Processing audio is one of the key functionalities provided by OpenAI. Developers can utilize advanced speech models (such as Whisper) to perform transformations, analysis, and generation of audio data, thereby supporting applications such as speech recognition and audio generation. For instance, developers might encounter random text in the output after the API recognizes audio [^40] https://community.openai.com/t/287544 or face InvalidRequestError when calling the whisper-1 API [^41] https://community.openai.com/t/433315, as shown in Fig. 9. Additionally, they often ask how to combine audio processing with image processing [^42] https://community.openai.com/t/22912. Challenges in this subcategory constitute 1.4% of all challenges.
Figure 9: A sample post in audio processing (C.5).
@@
Vision Capability (C.6). The Vision API uses pre-trained vision models to quickly and accurately identify objects, scenes, and features in images, providing developers with powerful image processing capabilities. This functionality is widely applicable in areas such as image recognition, intelligent surveillance, and medical image analysis. However, developers may encounter error messages when using the gpt-4-vision-preview API [^43] https://community.openai.com/t/546593 or find that it occasionally misinterprets information in images [^44] https://community.openai.com/t/504043. They also ask usage-related questions, such as how to add a parameter to view image resolution [^45] https://community.openai.com/t/477563. This subcategory constitutes 0.7% of all challenges.

### v3

@@ 4. RQ1: Popularity Trend Analysis
Fig. 3 illustrates the rising trend in LLM development popularity, as evidenced by the increasing number of posts on the OpenAI forum and the surge in registered users. The figure begins at the forum’s creation, tracking the growth in new posts and users every three months. Given that the user number significantly exceeds the post number, we apply a logarithmic transformation to the actual numbers for a clearer representation of the results. This visualization reveals a significant increase in developer interest in OpenAI since 2021, underscoring the timeliness and urgency of our empirical study on challenges for LLM application developers.
@@
Figure 3. The number of new posts and users every three months since creating the OpenAI developer forum. Notice, in the $y$-axis, we apply a logarithmic transformation to the actual numbers for a clearer representation of the results
Before November 2022, the growth trend was fairly consistent. However, the introduction of GPT-3.5 and its fine-tuned product, ChatGPT, to the public by OpenAI in late November 2022, marked a substantial shift. The number of posts increased by 292%, and the number of users rose by 1,302% compared to the preceding period. The release of GPT-4 to paying customers in March 2023 further accelerated this trend, with a 323% increase in posts and a 628% increase in users. This strong growth trend has persisted until now. Regarding the results of RQ1, we believe that this trend is not just good news for the LLM vendors (such as OpenAI), but has also driven the increasing adoption of LLMs within the software engineering community (Fan et al., 2023; Wang et al., 2024) and leads to the emergence of the LLM4SE (LLMs for Software Engineering) research field. The increasing use of LLMs suggests that their application is transforming software development, testing, and maintenance processes, especially in tasks such as code generation (Yan et al., 2023; Gu et al., 2024), code completion (Eghbali and Pradel, 2024), test generation (Schäfer et al., 2023; Yuan et al., 2023), program repair (Jin et al., 2023), vulnerability detection (Lu et al., 2024; Steenhoek et al., 2024), and source code summarization (Sun et al., 2024) (The latest research progress for these tasks can be referenced in Section 8.1). This trend highlights the significant role LLMs can play in enhancing productivity and efficiency across various stages of the software engineering lifecycle.
@@
Finding 1. The introduction of GPT-3.5 and ChatGPT in November 2022, followed by GPT-4 in March 2023, significantly accelerated the growth of the LLM developer community, markedly increasing the number of posts and users on the OpenAI developer forum.
5. RQ2: Difficulty Level Analysis
As described in Section 3, we evaluate the difficulty level of LLM-related questions from three different perspectives.
First, we use Fig. 4 to show the number of posts with different numbers of replies. The results show that 54% of questions receive fewer than three replies, while only 10% receive more than ten replies. This suggests that questions related to LLM development are generally challenging to resolve. However, we acknowledge that using the number of replies as a metric for assessing question difficulty has its limitations. For example, a greater number of responses to a question might suggest that reaching an agreement requires discussions from multiple developers or several iterations. On the other hand, questions with fewer replies could be simpler to solve, leading to less discussion.
Figure 4. The number of posts with different numbers of replies.
Then, we examine the proportion of posts with an accepted solution. Fig. 5 illustrates the percentage of posts that have an accepted solution. From these statistical results, we find that only 8.98% of the posts have accepted solutions. The relatively low acceptance rate suggests that these questions are generally more difficult to resolve. Moreover, a low acceptance rate may indicate that the quality assurance mechanism of the community’s posts is relatively immature compared to more established developer Q&A platforms like Stack Overflow, where accepted solutions are marked more frequently.
Figure 5. Proportion of posts with accepted solutions.
Finally, we analyze the time required to receive the first reply. This metric offers additional insight into the complexity of the questions, as a longer response time may indicate that the question is more difficult to address. Fig. 6 presents the average response time of questions with the replies for each subforum. As shown in this figure, most posts have a relatively long response time (approximately between 147 hours and 278 hours for different subforums), suggesting that many of these questions are challenging or that they require more time for developers to give a response.
Figure 6. The average time required to receive the first reply for each subforum.
@@ 5. RQ2: Difficulty Level Analysis
Based on the above results from three different perspectives, we summarize potential reasons as follows. First, the new and complex technological domains of LLM require deep expertise and skills from developers. This scarcity of qualified professionals results in limited responses. Second, the rapid evolution of LLM technology often leaves issues unresolved, leading to fewer available answers. Finally, the lack of comprehensive documentation and tailored support resources makes it difficult to address diverse developer needs, prolonging the resolution process for many questions.
@@
Finding 2. Problems related to LLM development often face challenges in receiving immediate, sufficient, and even accepted replies. This highlights potential directions for the community and LLM vendors to offer more effective support.
6. RQ3: Challenge Taxonomy Construction
@@
Fig. 7 illustrates the hierarchical taxonomy of challenges faced by LLM developers. In this figure, the nodes are shaded in descending levels of grey based on their depth in the hierarchy (notice leaf nodes are white). Each leaf node represents a specific category, while its parent node is an inner node composed of multiple subcategories. For example, API (B) is an inner category that can be further classified into three leaf categories: Faults in API (B.1), Error Messages in API Calling (B.2), and API Usage (B.3).
@@
The proportion of posts for each category is shown in parentheses. In summary, our taxonomy includes six inner categories and 26 leaf categories. Based on this taxonomy, we find that LLM developers encounter a wide variety of problems across different aspects of development, highlighting the diversity of challenges in LLM development. Next, we describe and provide examples for each category and summarize our findings and implications for developers and LLM vendors.
Figure 7. Our Constructed Challenge Taxonomy for LLM Developers.
Finding 3. The challenges faced by LLM developers are multifaceted and diverse, encompassing 6 categories and 26 distinct subcategories.
6.1. General Questions
@@
General questions encompass challenges that lack specific implementation details and are often posed by developers seeking foundational knowledge in LLM development. Our findings indicate that 26.3% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@
Integration with Custom Applications (A.1). With the launch of ChatGPT and the availability of LLM’s APIs, more developers are leveraging these GPTs/APIs to build custom applications, such as translation systems and text classification systems. Similar to the challenges in ML application integration (Alshangiti et al., 2019), we find that developers integrating LLMs with their applications also encounter difficulties, such as API configuration, environment setup, and the outputs of LLMs. For example, developers seek to understand how to integrate custom GPTs with web page interfaces [^9] https://community.openai.com/t/589076 or inquire about the feasibility of implementing specific functionalities through API calls, such as considering specialized terminology that varies across different languages [^10] https://community.openai.com/t/584762, as shown in Fig. 8. Additionally, some developers raise concerns about the reproducibility of results in commercial text classification systems using GPT-3.5-turbo, especially after model updates [^11] https://community.openai.com/t/281658. Developers also encounter problems related to hallucinations [^12] https://community.openai.com/t/408275, as shown in Fig. 9. Our findings indicate that challenges in this subcategory account for a significant 17.0%, highlighting the diverse and complex technical issues developers face when building their custom applications.
@@
Figure 8. A sample post in the integration with custom applications (A.1) related to specialized terminology.
Figure 9. A sample post in the integration with custom applications (A.1) related to hallucinations.
@@
Conceptional Questions (A.2). This subcategory encompasses questions about basic concepts or background knowledge of LLM development, such as how to calculate the tokens needed for each question [^13] https://community.openai.com/t/81018, whether the GPT-4 API has access to the Internet [^14] https://community.openai.com/t/468615, or whether plugins can be used with the API [^15] https://community.openai.com/t/186333. Additionally, developers inquire about the internal implementation details, such as mechanisms of the content filter [^16] https://community.openai.com/t/1095. In our taxonomy, the posts in this subcategory account for 6.4%.
@@
Finding 4. The majority (26.3%) of the total challenges fall under the category of General Questions. Within this category, the subcategory Integration with Custom Applications represents the largest proportion, accounting for 64.6% of General Questions.
@@
Discussion and Implication: General questions often reflect fundamental challenges developers face when working with LLM technologies. These challenges span various aspects (such as application integration, conceptual questions, and feature suggestions). Similar to the empirical studies (Zhang et al., 2019) on common challenges in developing deep learning applications, when leveraging APIs provided by OpenAI to build their applications (such as translation systems, text analysis systems, speech recognition systems, and others), developers often encounter complex technical questions. These include API integration methods, performance issues, output reproducibility issues, interpretability of output content, and so on. These challenges highlight the diverse and intricate nature of API integration and the need for clearer guidelines and examples from the LLM vendors to assist developers in these areas. As integration with custom applications is a major challenge, The LLM vendors could develop dedicated resources and support mechanisms to streamline this process. This may include detailed tutorials and integration best practices. Questions in the subcategory of conceptual questions involve basic concepts and background knowledge necessary for understanding OpenAI’s technologies. Developers often inquire about token calculations, API capabilities and internal mechanisms, model versions, and other related issues. These challenges underscore the need for improved foundational documentation. To mitigate these challenges, as suggested by Chen et al. (Chen et al., 2020), improving the usability and completeness of the documentation will help developers grasp fundamental concepts as their required skills more effectively, reducing the frequency of basic questions and accelerating their development process. For the last subcategory, developers frequently suggest new features or improvements related to API operations, such as better cost management tools, enhanced plugin and GPTs recommendation features, and support for a wide range of data formats. This subcategory reflects developers’ proactive interest in enhancing OpenAI’s tools. Therefore, the LLM vendors (e.g., the OpenAI organization) should establish a structured feedback mechanism to capture and address feature suggestions from developers. This will help prioritize feature development based on actual user needs and enhance overall user satisfaction.
6.2. API
@@
This category contains challenges related to the LLM APIs. In our taxonomy, 22.9% of the total challenges fall under this category. The detailed subcategories are discussed as follows.
@@
Faults in API (B.1). When calling LLM APIs, developers frequently encounter a variety of issues, such as low-quality generated content, limitations in model comprehension, and text coherence problems. These issues often result in LLM outcomes that do not meet developers’ expectations. The majority of these issues are related to unsatisfactory output, such as the presence of extraneous information (like spaces and newlines) in the API’s responses [^19] https://community.openai.com/t/578701, as well as phrase repetition in answers [^20] https://community.openai.com/t/54737. Beyond these common faults, developers face other types of faults, such as APIs failing to respond due to requests being too frequent [^21] https://community.openai.com/t/483090. Additionally, when dealing with specific technical problems, such as converting natural language into SQL queries, they might produce incorrect answers [^22] https://community.openai.com/t/426293. These examples illustrate the diverse technical challenges that can arise. This subcategory represents 8.7% of the total challenges.
@@
Error Messages in API Calling (B.2). This category primarily includes various error messages that developers encounter when calling LLM APIs, which is also faced by DL application developers (Morovati et al., 2024). For example, developers might face request errors [^23] https://community.openai.com/t/154711 and data value capacity limit errors when calling an API for image editing [^24] https://community.openai.com/t/464513. As shown in Fig. 10, developers may also face issues related to the OpenAI server, such as gateway timeout errors [^25] https://community.openai.com/t/219369. This subcategory represents 7.5% of the total challenges.
Figure 10. A sample post in the error messages in API calling (B.2).
@@
API Usage (B.3). Similar to challenges in developing deep learning applications (Zhang et al., 2019; Morovati et al., 2024), the parameter configuration is a principal concern among LLM developers, necessitating the parameter value adjustment to meet specific requirements and generate expected outputs from the API. A significant portion of these problems center on determining the appropriate values for API parameters [^26] https://community.openai.com/t/480243 (as shown in Fig. 11) and understanding the internal working and implementation mechanism (such as the response_model parameter [^27] https://community.openai.com/t/511738). Additionally, some challenges are closely associated with GPT models. For example, setting identical values for the seed parameter does not result in consistent outputs for specific GPT models [^28] https://community.openai.com/t/487245. This subcategory represents 6.7% of the total challenges.
Figure 11. A sample post in API usage (B.3).
@@
Discussion and Implication: The use of APIs is a challenge faced by developers in many fields, such as desktop applications (Scoccia et al., 2021), web applications (Samudio and LaToza, 2022), and deep learning applications (Cao et al., 2022; Morovati et al., 2024). Our results show that developers also frequently encounter various issues when using OpenAI APIs, leading to a decline in development efficiency. Some of these problems are related to the specific GPT models being used. For example, certain APIs only support the gpt-4 turbo model for specific fields like response_format, indicating potential compatibility issues. To mitigate confusion and reduce the frequency of these issues, the OpenAI organization should ensure that their documentation clearly outlines the compatibility of different models and fields, and provide detailed examples and best practices for developers to follow. Moreover, researchers can develop tools for detecting and repairing these compatibility issues by following the suggestions of Wang et al. (Wang et al., 2023) after analyzing compatibility issues in DL systems. Developers also encounter various error messages when calling OpenAI API. To address these issues, the OpenAI organization should consider providing more detailed error messages that help developers understand the cause of the errors and how to fix them. Additionally, offering guidance on common parameter configurations and their implications can be beneficial. For example, addressing phrase repetition issues can be managed by adjusting the frequency_penalty parameter, where setting positive values can penalize new tokens based on their existing frequency in the text so far, thereby decreasing the model’s likelihood of repeating the same line verbatim. Parameter configuration is a principal concern among developers. To help developers, the OpenAI organization should provide comprehensive documentation that includes detailed explanations of parameter values and their effects. Offering sample configurations and use-case scenarios can guide developers in making informed decisions about parameter settings. For example, explaining how different temperature parameter values affect the generated content’s creativity and randomness can help developers tailor the API output to their needs. In summary, LLM vendors should consider five important factors when designing API documentation (i.e., documentation of intent, code examples, matching APIs with scenarios, the penetrability of the API, and format and presentation), as suggested by Robillard and Deline (Robillard and DeLine, 2011) and improve API usability based on human-centered design methods (such as understanding how developers think about API functionality to design more intuitive interface, ensuring consistent parameter order and naming conventions, and avoiding long, confusing parameter lists), as suggested by Myers and Stylos (Myers and Stylos, 2016). Moreover, providing more informative error messages can help developers quickly identify and resolve issues. Finally, they should monitor community forums and address recurring issues by updating the API and documentation accordingly as suggested by Chen et al. when analyzing challenges in deploying deep learning-based software (Chen et al., 2020).
6.3. Generation and Understanding
@@
Due to powerful generation and understanding capabilities, developers can use LLMs to perform a variety of tasks (such as text processing via the LLM GPT-4o, image processing via the LLM DALL-E, and audio processing via the LLM Whisper) or fine-tune the OpenAI LLMs for downstream tasks. Notice the challenges in these tasks often overlap with other challenge categories, so a question (such as parameter setting issues related to image generation [^29] https://community.openai.com/t/359438) may be classified into multiple categories. Challenges related to this category account for 19.9% of the total challenges. The detailed subcategories are discussed as follows.
@@
Text Processing (C.1). As LLM’s core capability, text generation helps developers to create coherent and semantically sound textual content based on the input text (i.e., the prompt). This feature allows developers to automatically generate various types of text (such as articles, stories, and dialogues). By modifying the input prompts, developers can control the style, content, and length of the generated text to meet diverse application requirements across different scenarios. This category contains challenges associated with text generation, such as encountering errors when using the text generation API [^30] https://community.openai.com/t/486260 or getting repeated responses from the LLM davinci [^31] https://community.openai.com/t/254241. Moreover, developers want to provide feedback on the given output to improve the upcoming outputs when using the gpt-3.5-turbo model [^32] https://community.openai.com/t/452220. This subcategory constitutes 6.8% of all challenges.
@@
Fine-tuning GPT Models (C.2). LLM developers can fine-tune the models provided by the OpenAI organization by using the training datasets of the downstream tasks. Due to the widespread use of fine-tuning, developers frequently face questions related to this functionality. For example, they often inquire about how to set the parameter n_epochs [^33] https://community.openai.com/t/339246 or request the fine-tuned model to output results in JSON format [^34] https://community.openai.com/t/525169. Additionally, as shown in Fig. 12, developers are interested in whether they can train a fine-tuned model in an iterative way [^35] https://community.openai.com/t/18248. This subcategory of questions accounts for 6.7% of the total challenges.
Figure 12. A sample post in fine-tuning GPT models (C.2).
@@
Image Processing (C.3). Developers can influence image generation by inputting prompts into the model, allowing for the creation of images with customized features and styles. This versatile feature is useful across diverse applications, including image editing and artistic creation, offering developers opportunities for creative expression and design. However, developers often encounter challenges such as request errors when seeking image generation solutions [^36] https://community.openai.com/t/586006 or perceiving that the quality of the generated images is lower than expected [^37] https://community.openai.com/t/435187. Finally, they find that the image quality generated through API calls is inferior to that of the ChatGPT web version using the same prompts [^38] https://community.openai.com/t/435187. This subcategory constitutes 2.5% of all challenges.
@@
Embedding Generation (C.4). By using the Embedding APIs provided by LLM vendors, developers can leverage powerful text representation capabilities to convert text into high-dimensional vector representations. These vectors can capture the semantics and contextual information of the text, which can be used for tasks such as semantic search and text clustering. For example, developers may encounter PermissionError when creating embeddings [^39] https://community.openai.com/t/587579. Additionally, they often inquire whether modifying user queries can improve the accuracy of semantic search [^40] https://community.openai.com/t/393047 and how to determine the length of embedded content [^41] https://community.openai.com/t/111471. Challenges related to this subcategory account for 1.8% of the total challenges.
@@
Audio Processing (C.5). Processing audio is one of the key functionalities provided by the OpenAI organization. Developers can utilize advanced speech models (such as the LLM Whisper) to perform transformations, analysis, and generation of audio data, thereby supporting applications such as speech recognition and audio generation. For instance, developers might encounter random text in the output after the API recognizes audio [^42] https://community.openai.com/t/287544 or face InvalidRequestError when calling the LLM whisper API [^43] https://community.openai.com/t/433315, as shown in Fig. 13. Additionally, they often ask how to combine audio processing with image processing [^44] https://community.openai.com/t/22912. Challenges in this subcategory constitute 1.4% of all challenges.
Figure 13. A sample post in audio processing (C.5).
@@
Vision Capability (C.6). The Vision API uses pre-trained vision LLMs to quickly and accurately identify objects, scenes, and features in images, providing developers with powerful image processing capabilities. This functionality is widely applicable in areas (such as image recognition, intelligent surveillance, and medical image analysis). However, developers may encounter error messages when using the gpt-4-vision-preview API [^45] https://community.openai.com/t/546593 or find that it occasionally misinterprets information in images [^46] https://community.openai.com/t/504043. They also ask usage-related questions, such as how to add a parameter to view image resolution [^47] https://community.openai.com/t/477563. This subcategory constitutes 0.7% of all challenges.